# Non-Gaussianities in primordial black hole formation and induced gravitational waves

### v1

@@ V Induced Gravitational Waves
The curvature perturbation can source gravitational waves via nonlinear coupling terms in the perturbation theory up to second order. In this paper, we only briefly discuss the impact of the non-Gaussian curvature perturbation on the amplitude (via the normalization by the PBH abundance) as well as the shape of the induced GW spectrum. For a general discussion of scalar-induced GWs, see Witkowski et al. (2022); Domènech (2021) as well as Domènech (2024).
@@
where $\Phi=\frac{2}{3}\mathcal{R}$ is the curvature perturbation in longitudinal gauge calculated in the radiation dominated era, and $e^{ij}_{\mathbf{k}}$ is one of the polarization tensor. After solving the equation of motion (80), we can calculate the energy spectrum of the GWs generated at the horizon reentry, and evaluate it until today. Suppose the non-Gaussian curvature perturbation given by the local form (17), $\mathcal{R}=\mathscr{F}(\mathcal{R}_{\mathrm{g}})$, can be expanded as
@@
$\quad\cdot\left(\mathcal{P}_{\mathcal{R}}(uk)+F_{\text{NL}}^{2}\int^{\infty}_{0}\mathrm{d}\nu\int^{1+\nu}_{|1-\nu|}\mathrm{d}\mu~{}\frac{\mathcal{P}_{\mathcal{R}}(\mu uk)\mathcal{P}_{\mathcal{R}}(\nu uk)}{\mu^{2}\nu^{2}}+\cdots\right)$
@@
$\quad\cdot\left(\mathcal{P}_{\mathcal{R}}(vk)+F_{\text{NL}}^{2}\int^{\infty}_{0}\mathrm{d}\lambda\int^{1+\lambda}_{|1-\lambda|}\mathrm{d}\rho~{}\frac{\mathcal{P}_{\mathcal{R}}(\lambda vk)\mathcal{P}_{\mathcal{R}}(\lambda vk)}{\lambda^{2}\rho^{2}}+\cdots\right).$ (83)
Here we omit the contribution from the connected part $\langle\mathcal{R}^{4}\rangle_{c}$, which is negligibly small in most of the cases Adshead et al. (2021).
@@
The energy spectrum of the induced GW (83) can only be calculated for perturbative series like (81). Fortunately, because the induced GWs only depends on the power spectrum of the curvature perturbation $\mathcal{P_{R}}$ (at most $\sim 10^{-2}$ for enhanced peak), the higher order terms like $G_{\mathrm{NL}}\mathcal{R}^{3}$, $H_{\mathrm{NL}}\mathcal{R}^{4}$ contribute negligibly amount to the GW spectrum, as long as these coefficients $G_{\mathrm{NL}}$’s are not too large. Loose bounds for them are $G_{\mathrm{NL}}\ll\mathcal{O}(10^{2})F_{\mathrm{NL}}$, $H_{\mathrm{NL}}\ll\mathcal{O}(10^{4})F_{\mathrm{NL}}$, etc., which are easy to be satisfied for most of the models.
@@
There are two important examples. In the curvaton scenario with negligible curvaton at its decay, the perturbative series of (21) is valid even if $f_{\mathrm{NL}}\gg 1$, as the higher order terms ($G_{\mathrm{NL}}\mathcal{R}^{2}$ et al.) are always highly suppressed. In the ultra-slow-roll inflation with a sharp end, $\mathcal{R}(\mathcal{R}_{\mathrm{g}})$ is given by (60), of which the nonlinear parameters are
$F_{\mathrm{NL}}=3/2,\quad G_{\mathrm{NL}}=3,\quad H_{\mathrm{NL}}=27/4,\quad\cdots,$ (84)
which surely guarantees that higher order terms are suppressed. It was proved in Abe et al. (2022) that the contribution to the induced GW from the higher order terms of the logarithmic $\mathcal{R}(\mathcal{R}_{\mathrm{g}})$ are indeed negligible.
@@
Figure 7: Left: The spectrum of GWs induced by the enhanced narrow-peak power spectrum (76), with PBH as all the dark matter. The amplitude $\mathcal{A_{R}}$ for different parameters follows Table 1. From left to right, the black, purple dotted, blue dashed, cyan thick, brown dot-dashed, magenta dashed, and gray curves are for Gaussian case, $f_{\mathrm{NL}}=1$, $f_{\mathrm{NL}}=5/2$, ultra-slow-roll model, $f_{\mathrm{NL}}=10$, $f_{\mathrm{NL}}=10^{2}$, and $f_{\mathrm{NL}}=10^{3}$ respectively. The designed power-law integrated sensitivity curve of space-borne interferometers LISA (gray thick) Bartolo et al. (2016), Taiji (gray dashed) Wang and Han (2021), TianQin (gray dotted) Liang et al. (2021), and DECIGO (gray thin) Schmitz (2021) are also shown. Right: Two more realistic signals, the spectrum of GWs induced by ultra-slow-roll inflation with a sharp transition to slow-roll (blue), and a smooth transition to slow-roll, i.e. Starobinsky’s linear potential model (red). The right panel was drawn with the help of the SIGWfast Witkowski (2022) package.
@@
According to the current observational constraints, asteroid mass PBHs can serve as all the dark matter, of which the scalar induced GWs are right in the mHz band to be detected by the space-borne interferometers. So as an concrete example, we calculate the induced GW spectra for monochromatic power spectra (76) with different $f_{\mathrm{NL}}$’s, but normalized such that PBHs of $M\sim 2\times 10^{21}~{}\text{g}$ are all the dark matter. The resultant induced GWs peak in the millihertz band, which are shown in Figure 7 together with the designed sensitivity curves of the space-borne interferometers LISA Amaro-Seoane et al. (2013, 2012, 2017); Barausse et al. (2020), Taiji Ruan et al. (2020), TianQin Luo et al. (2016). The induced GW from Gaussian curvature perturbation with a monochromatic power spectrum was firstly derived numerically in Ananda et al. (2007) and analysically in Kohri and Terada (2018), which displays a resonant peak of logarithmic divergence at $(2/\sqrt{3})k_{*}$, and an infrared scaling of $\propto k^{2}$. As we commented before, positive $f_{\mathrm{NL}}$ can enhance the PBH formation, which requires less $\mathcal{A_{R}}$ for PBHs as all the dark matter, thus less GW spectrum. However, when $f_{\mathrm{NL}}\lesssim 10$, the suppression is smaller than one order of magnitude, which makes the shape of the GW spectrum around the peak similar to the Gaussian one and difficult to distinguish by the space-borne inteferometers. In this case, the GWs beyond the cutoff frequency $2k_{*}$ is purely from the non-Gaussian part, which is a smoking gun of primordial non-Gaussianity Cai et al. (2019a); Abe et al. (2022). Although these contributions are tiny compared with the main peak, they could be probed by the deci-hertz GW detector like BBO ($0.1$ to 1 Hz) Crowder and Cornish (2005); Corbin and Cornish (2006) or DECIGO ($10^{-2}$ to 1 Hz) Kawamura et al. (2006, 2011).
@@
As $f_{\mathrm{NL}}$ continues to increase, the main peak contributed by the Gaussian part decreases, yet the non-Gaussian contribution increases and changes the shape of the spectrum significantly. The non-Gaussian source is a convolution of two spectra with different wavenumbers, which gives a power-law function with a sharp UV cutoff. Therefore, the GW spectrum induced by the non-Gaussian part behaved as if it were induced by a power-law spectrum with a UV cutoff, which makes its infrared scaling $\Omega_{\mathrm{GW}}^{\mathrm{(NG)}}\propto k^{3}$, a property for a broad spectrum Cai et al. (2019b). Also, $\Omega_{\mathrm{GW}}^{\mathrm{(NG)}}$ has a mild bump instead of a sharp peak at $\sim\sqrt{3}k_{*}$, of which the amplitude is $\sim F_{\mathrm{NL}}^{4}\mathcal{A}_{\mathcal{R}}^{4}$. It becomes visible by LISA for $f_{\mathrm{NL}}\gtrsim 10$.
When $f_{\mathrm{NL}}$ becomes very large, the Gaussian part is invisible, while the non-Gaussian part dominates with an upper bound of $\Omega_{\mathrm{GW}}\sim 10^{-11}$, which is above the LISA sensitivity curve Cai et al. (2019a). Therefore, the detectability of scalar-induced GWs is more or less robust against the nonlinear parameter $f_{\mathrm{NL}}$, $\lambda_{-}$ (see (34) and (42)), or $r_{\mathrm{curvaton}}$ (see (51) and (52)), which made it an important scientific goal of LISA Barausse et al. (2020); Arun et al. (2022); Auclair et al. (2023); Bagui et al. (2023), Taiji Ren et al. (2023), and TianQin Liang et al. (2021) to probe the PBH dark matter.
@@
We should warn that the $\delta$-function peak we considered above is not physical. For more realistic models, the enhancement and decay of the power spectrum is milder, usually power law. Especially, in the typical model of ultra-slow-roll inflation, $\mathcal{P}_{\mathcal{R}}(k)$ increases as $\propto k^{4}$ before it reaches the maximum Byrnes et al. (2019). There might be modulated oscillations of period $\sim\pi$ if the transition from slow-roll to ultra-slow-roll is abrupt Starobinsky (1992); Biagetti et al. (2018); Cole et al. (2022); Pi and Wang (2023). The UV behavior of the power spectrum depends on the shape of the second slow-roll potential, usually displaying a power law $\mathcal{P_{R}}\propto k^{n_{s}-1}$. If the ultra-slow-roll stage ends sharply, the power spectrum drops off a little bit with new modulated oscillations with a larger period of $\sim\pi e^{N_{\mathrm{USR}}}$ Pi and Wang (2023). According to our discussion in Section III.3, these oscillations marks the sharp end of ultra-slow-roll stage, which are accompanied with the logarithmic relation (30). Based on the analytical expression of $\mathcal{P}_{\mathcal{R}}(k)$ as well as the statistical properties, we can draw the GW spectrum associated with a fixed PBH abundance. Currently, there is no satisfactory method of calculating PBH abundance for such a broad spectrum with non-Gaussianity, so we choose the same parameters in Table 1 to show schematically the induced GW spectrum in the ultra-slow-roll inflation in the right panel of Fig. 7. For comparison, we also draw the induced GWs from Starobinsky’s linear potential model, with represents the smooth end of ultra-slow-roll stage. We can see that similar to the $\delta$-function peak case, non-Gaussianity enhances the UV part of the induced GW spectrum, which makes the primary peak of the GW spectrum obscure. A drop off with modulated oscillations can be recognized as a smoking gun for the non-Gaussianity in the ultra-slow-roll inflation, which are detectable by DECIGO and BBO when PBHs constitute all the dark matter.
@@
When the transition from ultra-slow-roll to the second slow-roll stage is smooth, the curvature perturbation is dominated by the $\delta N$ contributed by the second slow-roll stage, which gives $f_{\mathrm{NL,2}}=(5/12)(n_{s,2}-1)$, where $n_{s,2}-1$ is the UV slope of $\mathcal{P_{R}}(k)$. In such induced GW spectra, there will be no modulated oscillations with period of $\sim\pi e^{N_{\mathrm{USR}}}$, and $f_{\mathrm{NL}}$ is determined by the UV slope of the GW spectrum, as $\Omega_{\mathrm{GW}}(k)\sim k^{2(n_{s,2}-1)}$. This can also be used to probe non-Gaussianity Atal and Domènech (2021).
@@
In summary, it is clearly shown in Figure 6 and Figure 7, the dependence of the induced GW on the non-Gaussianity is mild, once the PBH abundance is fixed. The impact of non-Gaussianity is mainly on the spectral shape, which could be detected in the future.

### v2

@@ V Induced Gravitational Waves
The curvature perturbation can source gravitational waves via nonlinear coupling terms like scale-scalar-tensor. In this paper, we only briefly discuss the impact of the non-Gaussian curvature perturbation on the amplitude (via the normalization by the PBH abundance) as well as the shape of the induced GW spectrum. For a general discussion of scalar-induced GWs, see Witkowski et al. (2022); Domènech (2021) as well as Chapter 18 Domènech (2024).
@@
where $\Phi=\frac{2}{3}\mathcal{R}$ is the curvature perturbation in longitudinal gauge calculated in the radiation dominated era, and $e^{ij}_{\mathbf{k}}$ is one of the polarization tensor. After solving the equation of motion (80), we can calculate the energy spectrum of the GWs generated at the horizon reentry, and evaluate it until today. For simple spectral shape of $\mathcal{R}$ like $\delta$-function peak Kohri and Terada (2018), lognormal peak Pi and Sasaki (2020), flat plateau Saito and Yokoyama (2010), etc., the GW spectrum can be calculated semi-analytically. Suppose the non-Gaussian curvature perturbation given by the local form (17), $\mathcal{R}=\mathscr{F}(\mathcal{R}_{\mathrm{g}})$, can be expanded as
@@
$\quad\cdot\left(\mathcal{P}_{\mathcal{R}\mathrm{g}}(uk)+F_{\text{NL}}^{2}\int^{\infty}_{0}\mathrm{d}\nu\int^{1+\nu}_{|1-\nu|}\mathrm{d}\mu~{}\frac{\mathcal{P}_{\mathcal{R}\mathrm{g}}(\mu uk)\mathcal{P}_{\mathcal{R}\mathrm{g}}(\nu uk)}{\mu^{2}\nu^{2}}+\cdots\right)$
@@
$\quad\cdot\left(\mathcal{P}_{\mathcal{R}\mathrm{g}}(vk)+F_{\text{NL}}^{2}\int^{\infty}_{0}\mathrm{d}\lambda\int^{1+\lambda}_{|1-\lambda|}\mathrm{d}\rho~{}\frac{\mathcal{P}_{\mathcal{R}\mathrm{g}}(\lambda vk)\mathcal{P}_{\mathcal{R}\mathrm{g}}(\lambda vk)}{\lambda^{2}\rho^{2}}+\cdots\right)+\cdots.$ (83)
Here for simplicity we omit the higher order terms as well as the contribution from the connected part $\langle\mathcal{R}^{4}\rangle_{c}$, which are negligibly small when $F_{\mathrm{NL}}\lesssim 10$. For the full expression, see Ref. Adshead et al. (2021) for quadratic expansion (up to $F_{\mathrm{NL}}$), Refs. Yuan et al. (2023); Li et al. (2023) for cubic expansion (up to $G_{\mathrm{NL}}$), Ref. Perna et al. (2024) for quintic expansion (81) (up to $I_{\mathrm{NL}}$), and Ref. Abe et al. (2022) for ultra-slow-roll inflation up to $\mathcal{A}_{\mathcal{R}}^{4}$.
@@
The energy spectrum of the induced GW (83) can only be calculated for perturbative series like (81). Fortunately, because the induced GWs mainly depends on the power spectrum of the curvature perturbation $\mathcal{P_{R}}$ which is at most $\sim 10^{-2}$ for an enhanced peak, the higher order terms like $G_{\mathrm{NL}}\mathcal{R}^{3}$, $H_{\mathrm{NL}}\mathcal{R}^{4}$ contribute negligibly to the GW spectrum, as long as these coefficients $G_{\mathrm{NL}}$’s are not too large. Loose bounds for them are $G_{\mathrm{NL}}\ll\mathcal{O}(10^{2})F_{\mathrm{NL}}$, $H_{\mathrm{NL}}\ll\mathcal{O}(10^{4})F_{\mathrm{NL}}$, etc., which are easy to be satisfied for most of the models.
@@
There are two important examples. In the curvaton scenario with negligible curvaton at its decay, the perturbative series of (21) is valid even if $f_{\mathrm{NL}}\gg 1$, as the higher order terms ($G_{\mathrm{NL}}\mathcal{R}^{2}$ etc.) are always strongly suppressed. In the ultra-slow-roll inflation with a sharp end, $\mathcal{R}(\mathcal{R}_{\mathrm{g}})$ is given by (57), of which the nonlinear parameters are
$F_{\mathrm{NL}}=3/2,\quad G_{\mathrm{NL}}=3,\quad H_{\mathrm{NL}}=27/4,\quad\cdots.$ (84)
This surely guarantees that higher-order contributions are negligible, which was shown explicitly up to $\mathcal{A}_{\mathcal{R}}^{4}$ in Ref. Abe et al. (2022).
@@
Figure 7: Left: The spectrum of GWs induced by the enhanced narrow-peak power spectrum (71). The amplitude $\mathcal{A_{R}}$ for different parameters follows Table 1, such that PBHs $\sim 10^{-12}M_{\odot}$ are all the dark matter, shown in Fig.6. From right to left, the orange, magenta dashed, brown dot-dashed, cyan thick, blue dashed, purple dotted, black thick, gray, and black dotted curves are for $f_{\mathrm{NL}}=10^{3}$, $f_{\mathrm{NL}}=10^{2}$, $f_{\mathrm{NL}}=10$, ultra-slow-roll, $f_{\mathrm{NL}}=5/2$, $f_{\mathrm{NL}}=1$, Gaussian case ($f_{\mathrm{NL}}=0$), curvaton scenario with $r\to 1$, and $f_{\mathrm{NL}}=-5/4$, respectively. The curves are drown by Adshead et al. (2023), following Adshead et al. (2021), together with the designed power-law integrated sensitivity curve of space-borne interferometers LISA (gray thick) Bartolo et al. (2016), Taiji (gray dashed) Wang and Han (2021), TianQin (gray dotted) Liang et al. (2021), and DECIGO (gray thin) Schmitz (2021). Right: Two more realistic signals, the spectrum of GWs induced by ultra-slow-roll inflation with a sharp transition to slow-roll (blue), and a smooth transition to slow-roll, i.e. Starobinsky’s linear potential model (red). The right panel was drawn by (83) with the help of the SIGWfast Witkowski (2022) package.
@@
According to the current observational constraints, asteroid mass PBHs can serve as all the dark matter, of which the scalar induced GWs are in the milliHz to deciHz band, to be explored by the space-borne interferometers. As an concrete example, we calculate the induced GW spectra for monochromatic power spectra (71) with different $f_{\mathrm{NL}}$’s, normalized such that PBHs of $M_{k_{*}}\approx 10^{-12}M_{\odot}$ are all the dark matter. The resultant induced GWs peak at $k_{*}=6.73\times 10^{-3}$ Hz, which are shown in Fig. 7 together with the power-law integrated sensitivity curves Thrane and Romano (2013) of the space-borne interferometers LISA Amaro-Seoane et al. (2013, 2012, 2017); Barausse et al. (2020), Taiji Ruan et al. (2020), TianQin Luo et al. (2016), and DECIGO Schmitz (2021). The induced GW from Gaussian curvature perturbation with a monochromatic power spectrum displays an infrared scaling of $\propto k^{2}$, a dip of zero at $\sqrt{2/3}k_{*}$, a resonant peak of logarithmic divergence at $(2/\sqrt{3})k_{*}$, and a sharp cutoff beyond $2k_{*}$ Kohri and Terada (2018). This infrared scaling is a consequence of the infinitesimal scale of the source Cai et al. (2019b), which is preserved even in the non-Gaussian case Adshead et al. (2021). As we commented before, positive $f_{\mathrm{NL}}$ can enhance the PBH formation, which requires less $\mathcal{A_{R}}$ for a fixed PBH abundance (e.g. $f_{\mathrm{PBH}}=1$), thus smaller GW spectrum is displayed. However, when $f_{\mathrm{NL}}\lesssim 10$, the suppression is within one order of magnitude, which makes the amplitude and shape of the GW spectrum similar to the Gaussian one and difficult to be distinguished. In this case, the GWs beyond the cutoff frequency $2k_{*}$ is purely from the non-Gaussian part, which is a smoking gun of primordial non-Gaussianity Cai et al. (2019a); Abe et al. (2022). Although these contributions are tiny compared with the main peak, they could be probed by the deci-hertz GW detector like BBO ($0.1$ to 1 Hz) Crowder and Cornish (2005); Corbin and Cornish (2006) or DECIGO ($10^{-2}$ to 1 Hz) Kawamura et al. (2006, 2011).
@@
As $f_{\mathrm{NL}}$ increases, the main peak contributed by the Gaussian part decreases due to the less requirement of $\mathcal{A_{R}}$ because of the enhanced productivity of PBHs. However, the non-Gaussian contribution increases slighly, and finally dominates the GW spectrum and changes its shape significantly. The non-Gaussian source is roughly a convolution of two spectra with different wavenumbers, which fills the dip at $\sqrt{2/3}k_{*}$. Also, $\Omega_{\mathrm{GW}}^{\mathrm{(NG)}}$ has a mild bump instead of a sharp peak at $\sim\sqrt{3}k_{*}$, of which the amplitude is $\sim F_{\mathrm{NL}}^{4}\mathcal{A}_{\mathcal{R}}^{4}$. It becomes visible by LISA for $f_{\mathrm{NL}}\gtrsim 20$.
A very important issue is that when PBHs constitute all the dark matter, $\Omega_{\mathrm{GW}}$ has a lower bound which is detectable by the space-borne interferometers Cai et al. (2019a). To see it clearly, note that the peak value of $\Omega_{\mathrm{GW}}$ is given by the maximum of the three peaks contributed by $\mathcal{O}(\mathcal{A}_{\mathcal{R}}^{2})$-, $\mathcal{O}(\mathcal{A}_{\mathcal{R}}^{3}f_{\mathrm{NL}}^{2})$-, and $\mathcal{O}(\mathcal{A}_{\mathcal{R}}^{4}f_{\mathrm{NL}}^{4})$-terms:
$\Omega_{\mathrm{GW,peak}}h^{2}$ $\approx 1.6\times 10^{-5}\mathrm{max}\Big{[}6.4\mathcal{A}_{\mathcal{R}}^{2},~{}3.7\mathcal{A}_{\mathcal{R}}^{3}F_{\mathrm{NL}}^{2},~{}3.9\mathcal{A}_{\mathcal{R}}^{4}F_{\mathrm{NL}}^{4}\Big{]},$
$\approx\mathrm{max}\left[\frac{4.5\times 10^{-9}}{(1+0.1f_{\mathrm{NL}})^{2}},~{}\frac{6.1\times 10^{-12}}{(1+0.1f_{\mathrm{NL}})^{3}}f_{\mathrm{NL}}^{2},~{}\frac{1.5\times 10^{-14}}{(1+0.1f_{\mathrm{NL}})^{4}}f_{\mathrm{NL}}^{4}\right].$
In the first expression, the numerical coefficient of the Gaussian contribution (“6.4”) comes from Ref. Pi and Sasaki (2020) with smoothing scale $\Delta=0.01$, while the coefficients of the $F_{\mathrm{NL}}^{2}$-term (“3.7”) and $F_{\mathrm{NL}}^{4}$-term (“3.9”) come from Ref. Adshead et al. (2021) (also with a width of $\Delta=0.01$). In the second equality, we use (79) for PBH as all the dark matter. From the above fitting formula, it is easy to see that $\Omega_{\mathrm{GW,peak}}h^{2}$ has a minimum of $\Omega_{\mathrm{GW,peak}}^{\text{(min)}}h^{2}\sim 8.4\times 10^{-11}$ when $f_{\mathrm{NL}}\sim 63$. After that, $\Omega_{\mathrm{GW}}$ only increase slightly, and finally approaches a constant of $1.1\times 10^{-10}$. This minimum is still well above the sensitivity curves of space-borne interferometers. Therefore, the detectability of scalar-induced GWs is robust against the nonlinear parameter $f_{\mathrm{NL}}$, $\lambda_{-}$ (see (34) and (42)), or $r$ (see (48) and (49)), which made it an important scientific goal of LISA Barausse et al. (2020); Arun et al. (2022); Auclair et al. (2023); Bagui et al. (2023), Taiji Ren et al. (2023), and TianQin Liang et al. (2021) to probe the PBH dark matter.
@@
We should warn that the $\delta$-function peak we considered above is not physical. For more realistic models, the enhancement and decay of the power spectrum around the peak is milder, usually obeying power-law. Especially, in the typical model of ultra-slow-roll inflation, $\mathcal{P}_{\mathcal{R}}(k)$ increases as $\propto k^{4}$ before it reaches the maximum Byrnes et al. (2019). There might be modulated oscillations of period $\sim\pi$ if the transition from slow-roll to ultra-slow-roll is abrupt Starobinsky (1992); Biagetti et al. (2018); Cole et al. (2022); Pi and Wang (2023). The UV behavior of the power spectrum depends on the shape of the second slow-roll potential, usually displaying another power-law $\mathcal{P_{R}}\propto k^{n_{s,2}-1}$. If the ultra-slow-roll stage ends sharply, the power spectrum drops off a little bit with new modulated oscillations with a larger period of $\sim\pi e^{N_{\mathrm{USR}}}$ Pi and Wang (2023). According to our discussion in Section III.3, these oscillations mark the sharp end of ultra-slow-roll stage, which are accompanied with the logarithmic relation (30). Based on the analytical expression of $\mathcal{P}_{\mathcal{R}}(k)$ as well as the statistical properties, we can draw the GW spectrum associated with a fixed PBH abundance. Currently, there is no satisfactory method of calculating PBH abundance for such a broad spectrum with non-Gaussianity, so we choose the same parameters as in the narrow peak case (marked “USR” in Table 1), aiming to demonstrate the feature but not the exact amplitude of the induced GW spectrum in the ultra-slow-roll inflation. This is shown in the right panel of Fig. 7. For comparison, we also draw the induced GW from Starobinsky’s linear potential model (with the narrow-peak amplitude), which represents the smooth end of ultra-slow-roll stage with negligible non-Gaussianity. We can see that similar to the $\delta$-function peak case, non-Gaussianity enhances the UV part of the induced GW spectrum, which makes the primary peak of the GW spectrum obscure. A drop off with modulated oscillations can be recognized as a smoking gun for the non-Gaussianity in the ultra-slow-roll inflation shown in (30), which are detectable by DECIGO and BBO when PBHs constitute all the dark matter.
@@
When the transition from ultra-slow-roll to the second slow-roll stage is smooth, the curvature perturbation is dominated by the $\delta N$ contributed by the second slow-roll stage, which has the normal consistency relation $f_{\mathrm{NL}}=(5/12)(1-n_{s,2})$ with $n_{s,2}$ the UV tilt of $\mathcal{P_{R}}(k)$. In such induced GW spectra, there will be no modulated oscillations with period of $\sim\pi e^{N_{\mathrm{USR}}}$, but $f_{\mathrm{NL}}$ is determined by the UV slope of the GW spectrum, as $\Omega_{\mathrm{GW}}(k)\sim k^{2(n_{s,2}-1)}$. This can also be used to probe non-Gaussianity Atal and Domènech (2021).
@@
In summary, it is clearly shown in Fig. 6 and Fig. 7 that the dependence of the induced GW amplitude on the non-Gaussianity is mild once the PBH abundance is fixed. However, the non-Gaussian impact on the spectral shape could be detected in the future.